Nutrition Calculations

1. A peanut-butter granola bar contains 2 grams of protein, 8 grams of fat, and 20 grams of carbohydrate. How many kcal in this bar come from protein?

2. One cheeseburger has 11 grams of carbohydrate and 340 kcal. What percentage of total kcal comes from carbohydrate? (round to nearest whole number)

3.  pouch of fruit snacks contains 19 grams of carbohydrate and 1 gram of protein. How many kcal are in one pouch?

4. A peanut-butter granola bar contains 2 grams of protein, 8 grams of fat, and 20 grams of carbohydrate. How many kcal in this bar come from carbohydrate?

5. A peanut-butter granola bar contains 2 grams of protein, 8 grams of fat, and 20 grams of carbohydrate. How many kcal in this bar come from fat?
Answer #1

1. A peanut butter granola bar contains 2 gms of protein = 2 \times 4 kcal/gm = 8 kcal

2. One cheeseburger has 11 grams of carbohydrate = 11 \times 4 kcal/gm = 44 kcal

Total calories = 340 kcal

The percentage of carbohydrate = 44 \div 340 \times 100 = 12.94% can be rounded to 13%

3. Pouch of fruit snacks contains 19 gms of carbohydrate = 19 \times 4 kcal/gm = 76 kcal

It contains 1 gm of protein = 1 \times 4 kcal/gm = 4 kcal

Total calories are 76 + 4 = 80 kcal

4. A peanut butter granola bar contains 20 grams of carbohydrate = 20 \times 4 kcal/gm = 80 kcal

5. A peanut butter granola bar contains 8 grams of fat = 8 \times 9 kcal/gm = 72 kcal